OpenStudy (anonymous):

simplify the radical expression 2√6 + 3√96

OpenStudy (anonymous):

simplify the radical expression (8+ √11)(8- √11)

OpenStudy (anonymous):

@jim_thompson5910 please????

jimthompson5910 (jim_thompson5910):

which one?

OpenStudy (anonymous):

both please?

jimthompson5910 (jim_thompson5910):

post one at a time please

jimthompson5910 (jim_thompson5910):

makes things a lot simpler

OpenStudy (anonymous):

simplify the radical expression 2√6 + 3√96

jimthompson5910 (jim_thompson5910):

anyways, did you get anything out of my hint that \[\Large \sqrt{96} = \sqrt{16*6}\] ??

jimthompson5910 (jim_thompson5910):

If not, then, \[\Large \sqrt{96} = \sqrt{16*6}\] \[\Large \sqrt{96} = \sqrt{16}*sqrt{6}\] \[\Large \sqrt{96} = 4*sqrt{6}\] Let me know if that helps or not.